1. 14.0 g sample of an aqueous solution of hydrobromic acid contains an unknown amount of the acid. If 25.6 mL of 0.517 M potassium hydroxide are required to neutralize the hydrobromic acid, what is the percent by mass of hydrobromic acid in the mixture?

__% by mass

2. A 13.1 g sample of an aqueous solution of nitric acid contains an unknown amount of the acid. If 18.7 mL of 1.18 M barium hydroxide are required to neutralize the nitric acid, what is the percent by mass of nitric acid in the mixture?

__% by mass
1
Expert's answer
2019-10-21T07:14:10-0400

1)

KOH + HBr = H2O +KBr

n(KOH) = 0.517*25.6/1000 = 0.0132 mol

n(KOH) = n(HBr) = 0.0132 mol

m( HBr) = 0.0132 * (1+81) = 1.0692 g

W( HBr) = 1.0692/14*100% = 7.64%


2)

2HNO3 + Ba(OH)2 = 2H2O + Ba(NO3)2

n(Ba(OH)2) = 1.18*18.7/1000 = 0.0221 mol

n(HNO3) = 2n(Ba(OH)2) = 0.0442 mol

m( HNO3) = 0.0442 * (1+14+48) = 1.39 g

W( HNO3) = 1.39/13.1*100% = 10.6%